# Multi-Grained Preference Enhanced Transformer for Multi-Behavior Sequential Recommendation

### v1

@@ Abstract.
Sequential recommendation (SR) aims to predict the next purchasing item according to users’ dynamic preference learned from their historical user-item interactions. To improve the performance of recommendation, learning dynamic heterogeneous cross-type behavior dependencies is indispensable for recommender system. However, there still exists some challenges in Multi-Behavior Sequential Recommendation (MBSR). On the one hand, existing methods only model heterogeneous multi-behavior dependencies at behavior-level or item-level, and modelling interaction-level dependencies is still a challenge. On the other hand, the dynamic multi-grained behavior-aware preference is hard to capture in interaction sequences, which reflects interaction-aware sequential pattern. To tackle these challenges, we propose a Multi-Grained Preference enhanced Transformer framework (M-GPT). First, M-GPT constructs a interaction-level graph of historical cross-typed interactions in a sequence. Then graph convolution is performed to derive interaction-level multi-behavior dependency representation repeatedly, in which the complex correlation between historical cross-typed interactions at specific orders can be well learned. Secondly, a novel multifaceted transformer architecture equipped with multi-grained user preference extraction is proposed to encode the interaction-aware sequential pattern enhanced by capturing temporal behavior-aware multi-grained preference . Experiments on the real-world datasets indicate that our method M-GPT consistently outperforms various state-of-the-art recommendation methods. Our code is available at: https://anonymous.4open.science/r/MGPT-DF31.
@@ 1. Introduction
  • CH2. Learning Dynamic Behavior-Aware Multi-Grained Preference.Sequential information is significant to multi-behavior sequential recommendation. A long-term interaction sequence can be divided into several sessions according to users’ dynamic multi-grained preference. As depicted in Figure 1, the interaction sequence of the boy consists of two main sessions. The first session including cell phones and earphones reflects the intention for electronic products. The second session including basketball shoes and basketball jerseys reveals the interest for sports. These two varying sessions show the dynamic user preference. From a deep perspective of a single session, when we only focus on one interaction, such as purchasing a pair of basketball shoes, we may only think that he is interested in collecting sneakers. However, when we combine other interactions, such as purchasing basketball jerseys, we will find that he is actually interested in playing basketball as a sport, which reveals the multi-grained preference. Meanwhile, we can distinguish the intensity of interest in varying items by different typed behaviors. Some previous works (e.g., MB-STR(Yuan et al., 2022), MBHT(Yang et al., 2022)) proposed methods to model the sequential information ignoring the dynamic behavior-aware multi-grained preference. Hence, how to model multi-behavior sequential pattern involving dynamic behavior-aware multi-grained preference is still a challenge for multi-behavior sequential recommendation.
@@
To solve aforementioned issues, in this paper, we propose Multi-Grained_Preference_enhanced_Transformer (M-GPT) to improve the recommendation performance. To achieve this goal, M-GPT is designed with two core components, e.g., interaction-aware dependency extractor and multifaceted sequential pattern generator. The interaction-level dependency extractor is developed to model the personalized interaction-level multi-behavior dependency from low-order to high-order(Ch1). In this component, we first construct a learnable graph structure according to item-behavior interactions in user historical sequence. Specifically, the incidence matrix of our graph is calculated by the inner product of item- and behavior-level dependency representations, which learns multi-behavior dependency at interaction-level. Then graph convolution is utilized iteratively to model interaction-level dependency in various orders. The multifaceted sequential pattern generator aims to capture the sequential interaction pattern enhanced by extracting behavior-aware multi-grained preference in different time scales (Ch2). We first perform the transformer layer with linear attention to model the sequential pattern more efficiently. Moreover, the interaction sequences are divided into several sessions by different time scales. Then the multi-grained self-attention is performed to capture session-specific multi-behavior preference based on multi-grained user intent in each session. To aggregate the multi-grained preference captured in each session, an aggregator is performed to fuse session-based multi-grained preference representations with temporal effects into our sequential interaction pattern. To summarize, the contributions in this paper are listed as follows,
@@ 4.3. Ablation Study
4.4. Study On Behavior Pattern
To investigate the ability of learning users’ diverse behavior pattern, we conduct comprehensive experiment on Taobao dataset. Specifically, we analyze specific sequences from the historical interactions of user89 and user2035. We consider all behaviors as the target behavior to observe the behavior patterns of each user. Figure 4 illustrates three 4x4 behavior-level dependency matrices learned by M-GPT. As shown in figure 5, we can observe the following: 1) Users’ behavior patterns vary according to their personal shopping habits. For instance, from figure 5 we can infer that user89 prefers to tag items as favorites before purchasing them. However, it appears that user2035 has no clear preference before making a purchase. 2) In the Taobao dataset, Page-view is the most important behavior. In figure 5 Taobao behavior pattern, Page-view has the highest relevance scores with all the other behaviors, which aligns with the common understanding that Page-view provides the first impression of items to users and triggers them to perform the next behavior.
4.5. Case Study
To offer an intuitive impression of our model interpretability, we conduct comprehensive case studies on Taobao, IJCAI and Retail dataset. We randomly select 100 items as candidate items for each datasets respectively. First, we study the attention map in case study. As figure 3 depicted, we show the multi-grained attention maps of two attention heads from two sequential sessions in Taobao and IJCAI datasets. It’s obvious that the multi-grained preference in each session can be well captured by various query in specific granularity. At last, we investigate the average scores on 100 candidate items from different multi-behavior dependency orders. We show the results in three datasets in figure 5. In conclusion, we find that learning multi-behavior dependency in different orders truly influences the recommendation performance. Specifically, for Taobao dataset, it’s more efficient to capture multi-behavior dependency at 2nd-order. Nevertheless, it seems that there is no notable gap in model performance at multi-order multi-behavior dependency for IJCAI and retail dataset.
@@ 5. Conclusion
In this paper, we propose a novel model M-GPT for multi-behavior sequential recommendation problem, which explores multi-order interaction-level dependency and multifaceted sequential pattern enhanced with multi-grained preference. The M-GPT consists of two components, i.e., interaction-level dependency extractor and multifaceted sequential pattern generator. In interaction-aware dependency extractor, we leverage graph-based method to learn the correlations among interactions from item-view and behavior-view information. Meanwhile, we utilize graph convolution to learn dependency representation in different levels. In multifaceted sequential pattern generator, we learn sequential pattern of user interaction sequence by linear self-attention mechanism and extract multi-grained user preference in different time scales to enrich the representation of sequential pattern. At last, we conduct comprehensive experiments on two public datasets, which further verifies the effectiveness of our proposed M-GPT compared with some state-of-the-art methods. In the future, we aim to merge the capability of GNN and LLM to achieve better recommendation performance.
Acknowledgements.
To Robert, for the bagels and explaining CMYK and color spaces.
@@ Appendix C Details of Baselines
(a) Mask Ratio
(b) Dependency order
(c) Number of Session
(d) Preference Granularity
(e) Mask Ratio
(f) Dependency order
(g) Number of Session
(h) Preference Granularity
(i) Mask Ratio
(j) Dependency order
(k) Number of Session
(l) Preference Granularity
[…]

### v2

@@ Abstract.
Sequential recommendation (SR) aims to predict the next purchasing item according to users’ dynamic preference learned from their historical user-item interactions. To improve the performance of recommendation, learning dynamic heterogeneous cross-type behavior dependencies is indispensable for recommender system. However, there still exists some challenges in Multi-Behavior Sequential Recommendation (MBSR). On the one hand, existing methods only model heterogeneous multi-behavior dependencies at behavior-level or item-level, and modeling interaction-level dependencies is still a challenge. On the other hand, the dynamic multi-grained behavior-aware preference is hard to capture in interaction sequences, which reflects interaction-aware sequential pattern. To tackle these challenges, we propose a Multi-Grained Preference enhanced Transformer framework (M-GPT). First, M-GPT constructs an interaction-level graph of historical cross-typed interactions in a sequence. Then graph convolution is performed to derive interaction-level multi-behavior dependency representation repeatedly, in which the complex correlation between historical cross-typed interactions at specific orders can be well learned. Secondly, a novel multifaceted transformer architecture equipped with multi-grained user preference extraction is proposed to encode the interaction-aware sequential pattern enhanced by capturing temporal behavior-aware multi-grained preference . Experiments on the real-world datasets indicate that our method M-GPT consistently outperforms various state-of-the-art recommendation methods. Our code is available at: https://anonymous.4open.science/r/MGPT-DF31.
@@ 1. Introduction
  • CH2. Learning Dynamic Behavior-Aware Multi-Grained Preference.Sequential information is significant to multi-behavior sequential recommendation. A long-term interaction sequence can be divided into several sessions according to users’ dynamic multi-grained preference. As depicted in Figure 1, the interaction sequence of the boy consists of two main sessions. The first session including cell phones and earphones reflects the intention for electronic products. The second session including basketball shoes and basketball jerseys reveals the interest for sports. These two varying sessions show the dynamic user preference. From a deep perspective of a single session, when we only focus on one interaction, such as purchasing a pair of basketball shoes, we may only think that he is interested in collecting sneakers. However, when we combine other interactions, such as purchasing basketball jerseys, we will find that he is actually interested in playing basketball as a sport, which reveals the multi-grained preference. Meanwhile, we can distinguish the intensity of interest in varying items by different typed behaviors. Some previous works (e.g., MB-STR(Yuan et al., 2022), MBHT(Yang et al., 2022)) proposed methods to model the sequential information ignoring the dynamic behavior-aware multi-grained preference. MISSL(Wu et al., 2024) emphasizes on capturing multi-typed interests. Hence, how to model multi-behavior sequential pattern involving dynamic behavior-aware multi-grained preference is still a challenge for multi-behavior sequential recommendation.
@@
To solve aforementioned issues, in this paper, we propose Multi-Grained Preference enhanced Transformer (M-GPT) to improve the recommendation performance. To achieve this goal, M-GPT is designed with two core components, e.g., interaction-aware dependency extractor and multifaceted sequential pattern generator. The interaction-level dependency extractor is developed to model the personalized interaction-level multi-behavior dependency from low-order to high-order(Ch1). In this component, we first construct a learnable graph structure according to item-behavior interactions in user historical sequence. Specifically, the incidence matrix of our graph is calculated by the inner product of item- and behavior-level dependency representations, which learns multi-behavior dependency at interaction-level. Then graph convolution is utilized iteratively to model interaction-level dependency in various orders. The multifaceted sequential pattern generator aims to capture the sequential interaction pattern enhanced by extracting behavior-aware multi-grained preference in different time scales (Ch2). We first perform the transformer layer with linear attention to model the sequential pattern more efficiently. Moreover, the interaction sequences are divided into several sessions by different time scales. Then the multi-grained self-attention is performed to capture session-specific multi-behavior preference based on multi-grained user intent in each session. To aggregate the multi-grained preference captured in each session, an aggregator is performed to fuse session-based multi-grained preference representations with temporal effects into our sequential interaction pattern. To summarize, the contributions in this paper are listed as follows,
@@ 4.3. Ablation Study
4.4. Behavioral Sequential Pattern Analysis
To investigate the ability of learning users’ diverse behavioral sequential pattern, we conduct comprehensive experiment on Taobao dataset. Specifically, we analyze specific sequences from the historical interactions of user89 and user2035. Figure 4 illustrates three 4x4 matrices calculated by the average of attention weights in global sequential pattern. Specifically, we treat behavior pairs and their reversed order (e.g. click-purchase and purchase-click) as different patterns considering various sequential encoding. As shown in figure 5, we can observe the following: 1) Users’ behavior patterns vary according to their personal shopping habits. For instance, from figure 5 we can infer that user89 prefers to tag items as favorites before purchasing them. However, it appears that user2035 has no clear preference before making a purchase. 2) In the Taobao dataset, Page-view is the most important behavior. In figure 5 Taobao behavior pattern, Page-view has the highest relevance scores with all the other behaviors, which aligns with the common understanding that Page-view provides the first impression of items to users and triggers them to perform the next behavior.
4.5. Exploratory Case Study
To offer an intuitive impression of our model interpretability, we conduct comprehensive case studies on Taobao, IJCAI and Retail dataset. We randomly select 100 items as candidate items for each datasets respectively. First, we study the attention map in case study. As figure 3 depicted, we show the multi-grained attention maps of two attention heads from two sequential sessions in Taobao and IJCAI datasets. It’s obvious that the multi-grained preference in each session can be well captured by various query in specific granularity. At last, we investigate the average scores on 100 candidate items from different multi-behavior dependency orders. We show the results in three datasets in figure 5. In conclusion, we find that learning multi-behavior dependency in different orders truly influences the recommendation performance. Specifically, for Taobao dataset, it’s more efficient to capture multi-behavior dependency at 2nd-order. Nevertheless, it seems that there is no notable gap in model performance at multi-order multi-behavior dependency for IJCAI and retail dataset. Moreover, we perform exploratory case study on mulit-behavior dependency in Appendix E.
@@ 5. Conclusion
In this paper, we propose a novel model M-GPT for multi-behavior sequential recommendation problem. In interaction-level dependency extractor, we leverage graph-based method to learn the correlations among interactions from item-view and behavior-view information. In multifaceted sequential pattern generator, we learn sequential pattern of user interaction sequence by linear self-attention mechanism and extract multi-grained user preference in different time scales to enrich the representation of sequential pattern. At last, we conduct comprehensive experiments on two public datasets verifying the effectiveness of our proposed M-GPT compared with some state-of-the-art methods. In the future, we aim to explore the capability of Graph Foundation Model (GFM) in Recommendation.
@@ Appendix B Time Complexity Analysis
(a) Mask Ratio
(b) Dependency order
(c) Number of Session
(d) Preference Granularity
(e) Mask Ratio
(f) Dependency order
(g) Number of Session
(h) Preference Granularity
(i) Mask Ratio
(j) Dependency order
(k) Number of Session
(l) Preference Granularity
[…]
@@ Appendix C Details of Baselines
  • NextIP(Luo et al., 2022). A Transformer based model that divides the problem into two subtasks: next-item prediction and purchase prediction
@@
  • DyMuS+(Cho et al., 2023). An enhanced version of DyMuS, where the dynamic GRU constructs its internal hidden states as capsules to further capture item-level correlations.
  • TGT(Xia et al., 2022). A multi-behavior SRS that captures short-term user interests with a behavior-aware transformer network and long-term user interests via a temporal graph neural network.
  • PBAT(Su et al., 2023). A transformer-based model that explores personalized multi-behavior patterns and multifaceted time-evolving collaborations for MBSR problem.
  • MISSL(Wu et al., 2024). It proposes a multi-interest self-supervised learning including a behavior-aware multi-interest encoder and intra- and inter-interest self-supervised learning.
  • END4Rec(Han et al., 2024). It efficiently captures intricate patterns in user behavior by eliminating noise from user behavior data through noise-decoupling contrastive learning and a guided training strategy.
@@ Appendix D Hyper-parameter Analysis
(a) MBHT, IJCAI
(b) M-GPT, IJCAI
(c) MBHT, Taobao
(d) M-GPT, Taobao
Figure 7. study on behavioral relationship
[…]
(a) incidence matrix of user1
(b) incidence matrix of user23
Figure 8. case study on interaction-level multi-behavior dependency
[…]
Appendix E MULTI-BEHAVIOR DEPENDENCY ANALYSIS
E.1. Behavioral Relationship
We present the $4\times 4$ behavioral relationship matrices for MBHT and M-GPT across two datasets, calculated using the $BB^{T}$ method. Based on these results, several key observations can be made: 1) The relational values between each pair of behaviors in M-GPT are significantly higher than those observed in MBHT. This underscores M-GPT’s advantage in explicitly modeling dependencies at the behavior level, capturing multi-behavior interactions more effectively. 2) Specifically, the relational values for ’favorite-purchase’ and ’add to cart-purchase’ pairs in M-GPT are markedly higher compared to MBHT. Given that users are generally more inclined to purchase items they have favorited or added to their carts, these findings provide insight into the superior performance of M-GPT in reflecting real-world user behavior patterns. 3) In the IJCAI dataset, the relational scores for ’add to cart’ interactions are notably low. This phenomenon could be attributed to the relatively sparse occurrence of ’add to cart’ events within this dataset, suggesting that the frequency of specific behaviors may influence the strength of their detected relationships.
In summary, the analysis of the behavioral relationship matrices highlights M-GPT’s capability to better capture complex interdependencies among user behaviors. Furthermore, the lower relational scores for certain behaviors, such as ’add to cart’, emphasize the importance of considering behavior frequency when interpreting model outcomes.
E.2. Interaction-Level Dependency
We also present the incidence matrices of our interaction-level graph in Figure 8. These matrices encapsulate the values of interaction-level dependencies within users’ historical sequences, calculated using equations (3)–(5). Several interesting observations can be drawn from these figures: 1) Both users exhibit strong long-term dependency interactions over time. 2) Short-term strong interaction-level dependencies are evident in both users’ sequences, likely attributable to temporary deviations in user preferences. 3) Notably, User 23 demonstrates a more stable and stronger multi-behavior dependency at the interaction level compared to User 1, reflecting distinct shopping habits between the two users.
These findings highlight the nuanced patterns of interaction-level dependencies, offering insights into the temporal dynamics and stability of user behaviors. The differences observed between the two users underscore the importance of considering individual behavioral characteristics when analyzing interaction data.